OpenStudy (anonymous):

A pile of gravel is shaped like a cone and has a radius of 6 ft and a height of 9 ft. Garden Landscaping Company purchases some of this gravel. After the gravel that Garden Landscaping Company purchases is removed, the pile has a radius of 3 ft and a height of 6 ft. How much gravel did Garden Landscaping Company purchase? Use 3.14 to approximate pi and express your final answer in tenths. @phi

OpenStudy (phi):

I guess we have to find the volume of the starting cone as a first step. can you do that ? I would use volume of a cone = 1/3 * pi *r^2 *h

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

v=1/3pir^2 v=1/3*3.14*(6)^2*(9) v=1/3*3.14*36*9 V=1/3*1017.36 v=1017.36/3 v=339.12

OpenStudy (phi):

ok now find the volume of the small cone

OpenStudy (anonymous):

v=1/3pir^2h v=1/3*3.14*(3)^2*(6) v=1/3*3.14*9*6 v=1/3*169.56 v=169.56/3 v=56.52

OpenStudy (phi):

ok, they started with the big amount 339.12, and took some away. what was left was 56.52 how much did they take away ?

OpenStudy (anonymous):

okay do i have to subtract those two

OpenStudy (phi):

if it is not clear, think about a smaller problem. if you started with 10 marbles, and your friend took some, and now you have 1 marble, how many marbles did your friend take ?

OpenStudy (anonymous):

oh 9 marbles

OpenStudy (phi):

and how do you get 9 ? 10 - 1 big - small use that for your problem

OpenStudy (anonymous):

i would then get 282.6 but thing is i have to express my answer in the hundreths place

OpenStudy (phi):

it says in your question express your final answer in tenths to be accurate to 100ths place, you should use a more exact value of pi here we used pi = 3.14 (notice that is one more digit than 10ths) and at the end, we round the answer to the nearest 10th to get 100ths, we would use at least 3.142 (1 more digit than 100ths) then round at the end. better is to use 3.14159, and then round at the end. but, any way, your answer is ok as it is for the nearest 10th

OpenStudy (anonymous):

oh okay thanks